INAF queue observing, 2016 Mar 01 - Mar 02

Observer: R. Carini, M. Faccini
Telescope Operator: David Gonzalo-Huerta
Support Astronomer: D Thompson on call
Instrument: LUCI, LBC

Summary

At the sunset the sky is covered, wind velocity around 5 m/s, humidity 40%, Temperature 4 C.

Night duration (12\xB0 el twilight) 19:30-5:30 ( 02:30-12:30 UT) = 10 h 00m

Observing time: 8h m (80%)

Weather time loss: * h* ()

technical time loss 2h m (20%)

Data Summary

Project_RUN Objects Filter ExpTime Notes
Antoniucci V1118 Ori 200 H+K grating, zJspec, HKspec 0.6 h completed
telluric star HIP33723 200 H+K grating, zJspec, HKspec    
Raimondo Uma II LBC-b: S,B; LBC-r: r 1 h 13 m two images of the field 8 have to be taken
Marconi IZw18 LBC-b: V; LBC-r: I 10m COMPLETED
Phot std S101 LBC-b: S,V,B,g,r; LBC-r: V, R, r, i, z    
deBarros M103GLASS2c 200 H+K grating, zJspec 1h 30m  
telluric      
flat and arc LUCI Antoniucci zJ and HKspec    

Details

02:30 the weather is improved, but the sky is completely veiled, we begin with LUCI2
  • begin with the Antoniucci program
  • 02:53 the TO has problem in the collimation due to the shell.
  • the TO tries again the collimation but he fails
  • 03:03 the TO try again, the problem could be to to the setting of the adaptive on the secondary mirror, but it should be static.
  • 03:10 the problem is not resolved
  • 03:50 problem resolved
  • 03:51 load the script for the target V1118 for the Antoniucci's progrma
  • 03:52 lost collimation again
  • 04:05 seems that the problem is due to GCS

04:08 GO the script v1118_ori.xml
  • acquisition luci2.20160302.0002-0003
  • 04:22 lost the guiding again

04:40 re-start science. GO v1118_ori.xml
  • acquisition luci2.20160302.0002-0008
  • seeing 0.8''
  • science: luci2.0009-0016

05:06 telluric star GO v1118_ori_tell.xml
  • acquisition luci2.0017-0023
  • science luci2.0024-0025

05:24 the sky is clear. we switch to LBC *05:44 problems with the connection of LBC
  • line of error: "Red channel: one or more system failed: FILTERS has problems on power line"

06:02 problem resolved. slew to Raimondo's program, field 8
  • run copointing UMaII _f8_UBr_point.ob
  • dofpia
  • lbcrangebal
  • 06:06 run science UMaII _f8_UBr.ob
  • seeing 0.8''- 0.9''
  • the first image in the red channel maybe is not so good
  • 07.09 seeing 0.75''' * the blue channel is stucked, the last two images are not taken

07:15 TO restart LBC

07:33 slew to field 9 of the Raimondo's program
  • dofpia
  • 07:37 run run science
  • seeing about 0.9'''
  • 08:09 seeing 1.15
  • 08:15 airmass 1.306 seeing 1.0'''

8:28 switch to Marconi's program
  • run IZw18_copointing.ob
  • dofpia
  • lbcrangebal
  • 08.38 RB_Science and iraf have lost the connection. We restart the programs
  • 08:42 run the science IZw18.ob
  • 08:50 seeing about 0.7'''

08:56 slew to the photometric standard S101

09:12 switch to LUCI2

09:31 slew to deBarros program
  • 09:41 again problems with the collimation

10:11 problems resolved, we can load the script M103GLASS2c.xml
  • acquisition luci2.20160302.0026-0031
  • seeing ~ 0.7''-0.8''
  • 10:42 go science luci2.0032-0037

11:16 GO M103GLASS2c.xml luci2.0038-0043
  • seeing ~0.8''- 0.9'''

11:51 GO M103GLASS2c.xml luci2.0044-0049
  • seeing ~0.8''-0.9''
  • 12:23 seeing 1-1.1

12:26 slew to telluric star
  • GO M103GLASS2c _tell.xml
  • acquisition luci2.0050-0054
  • science luci2.0055-0062=

13:05 the dome is close

13:10 begin the calibration
  • flat and arc for Antoniucci
Attention the flats are done manually, we have not changed the name of the object, so in the header of these data the object is the name of the telluric 'HIP59861'.

-- %USERSIG{RobertaCarini - 2016-03-02}%

Comments

 
Topic revision: r7 - 03 Mar 2016, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback